# Unifying Demonstration Selection and Compression for In-Context Learning

### v2

@@ Abstract
In-context learning (ICL) enables LLMs to exhibit spectacular emergent capabilities in various scenarios. Unfortunately, introducing demonstrations easily makes the prompt length explode, bringing a significant burden to hardware. In addition, random demonstrations usually achieve limited improvements in ICL, necessitating demonstration selection among accessible candidates. Previous studies introduce extra modules to perform demonstration compression or selection independently. In this paper, we propose an ICL framework UniICL, which Unifies demonstration selection and compression, and final response generation via a single frozen LLM. UniICL leverages the understanding ability of well-trained LLMs to independently compress different demonstrations into compressed features, and then a learnable projection layer converts features to LLM-acceptable compressed virtual tokens. Apart from substituting original demonstrations to reduce input length, virtual tokens are again used to select potential demonstrations. Finally, current queries together with selected compressed virtual tokens are fed into the same frozen LLM for response generation. UniICL is a parameter-efficient framework that only contains 17M trainable parameters originating from the projection layer and a learnable embedding. We build UniICL upon two backbones and conduct experiments over in- and out-domain datasets of both generative and understanding tasks, encompassing ICL scenarios with plentiful and limited demonstration candidates. Results show that UniICL effectively unifies 12×12$\times$12 × compression, demonstration selection, and response generation, efficiently scaling up the baseline from 4-shot to 64-shot ICL with 24 GB CUDA allocation [^1] The code and model will be released in the final version..
Unifying Demonstration Selection and Compression for In-Context Learning
Jun Gao1, Ziqiang Cao1∗, Wenjie Li2 School of Computer Science and Technology1 Department of Computer Science, Hong Kong Polytechnic University2
Figure 1: UniICL compresses candidate demonstrations into spans of compressed virtual tokens and selects demonstrations similar to the query. Virtual tokens substitute the original demonstration together with the query, fed to the same LLM for prediction generation.
@@ 1 Introduction
In-context learning (ICL) Brown et al. (2020); Xie et al. (2021); Wang et al. (2023b) exhibits powerful performance in practical applications with the emergence of scaled-up Transformer-based Vaswani et al. (2017) Large Language Models (LLMs) Wang et al. (2023c); Yang et al. (2023); Wei et al. (2023); Wang et al. (2023a); Min et al. (2022). In ICL, the provided demonstrations activate the pre-training knowledge of LLMs to allow them to perform well on various downstream tasks such as text summarization Wang et al. (2023c); Yang et al. (2023) and text classification Min et al. (2022) without gradient updating of billion parameters. Despite its significant role in the era of LLMs, ICL also brings an enormous challenge to the input window. Specifically, inevitably introducing demonstrations directly causes length explosion Wang et al. (2024), bringing significant memory costs and decreasing inference throughput as described in Figure 8. Except for length explosion, Liu et al. (2021) points out that the quality of the selected demonstrations significantly influences the ICL performance.
Recent efforts in modifying model architecture significantly expand the input window Zheng et al. (2022); Wu et al. (2022); Ding et al. (2023); Bulatov et al. (2023). However, LLMs with million context windows still struggle to overcome performance degradation Liu et al. (2024). Researchers also attempted to alleviate length explosion via prompt pruning Jiang et al. (2023) or soft prompts Wingate et al. (2022); Mu et al. (2023); Ge et al. (2023). Their main idea is to train an independent compressor to compress the input into soft prompts. While in selecting powerful in-context demonstrations, Liu et al. (2021); Lu et al. (2021) scores each candidate demonstration for the current queries via an extra ranker, and Ram et al. (2023); Wang et al. (2024) finds that a fine-tuned LLM is up to the task of ranking.
However, the standalone compressor or ranker inevitably incurs additional memory costs as it necessitates simultaneous loading alongside the target LLM. To tackle these challenges, we propose a framework UniICL as illustrated in Figure 1, which leverages the semantic understanding ability of target LLMs developed during pre-training to compress and select demonstrations. UniICL keeps the target LLM frozen during training to avoid catastrophic forgetting and reduce training costs. Considering the frozen LLM within UniICL function as both compressor and generator, UniICL is training efficiently and does not need to load extra compression modules, significantly reducing memory costs for both training and inference. As the underlying LLMs in visual language models fail to understand visual features without an adapter, UniICL still requires converting compressed features into LLM-acceptable compressed virtual tokens. Apart from substituting lengthy demonstrations, virtual tokens are dense representations of original demonstrations, which can be naturally further applied to select potential demonstrations by measuring similarity. In this paper, the connector is a projection layer supervised-tuned under the language modeling objective of the generator during learning compression and jointly optimized by language modeling and contrastive objectives He et al. (2020) during learning selection.
UniICL notices the fact that demonstrations in the ICL are independent of each other, which is ignored by previous compression studies Mu et al. (2023); Wingate et al. (2022); Ge et al. (2023). Therefore, UniICL independently compresses demonstrations and proposes to configure Demonstration Bank (DB) to cache compressed virtual tokens for further reusing without requiring repeated compression for the same demonstration. We evaluate UniICL which only contains 17M trainable parameters on a scope of benchmarks involving linguistic acceptability, semantic classification, text summarization, and passage reranking, and UniICL achieves outstanding performances. Then, the in-domain experiments show that UniICL built upon two backbones (Vicuna and BlueLM) effectively substitutes the 12$\times$ demonstrations with soft prompts, easily scaling up the inner LLM from 4-shot to 64-shot. Our main contributions are as follows:
  • To our knowledge, we are the first to propose an ICL framework that unifies compression, selection, and generation via a single frozen LLM.
  • UniICL is a memory-friend framework that enables LLMs to perform large-shot ICL on consuming GPUs.
@@
  • UniICL proposes to configure Demonstration Bank to enhance ICL efficiency, avoiding repeated compression for the same demonstration.
@@ 2.1 Soft Prompt Compression
Recently, researchers attempted to utilize soft prompts to convert actual tokens to dense-information virtual tokens. Mostly from a distillation perspective, Wingate et al. (2022) aligned the teacher model and the student model, where the teacher model accepted the actual task instruction while the student model fed the soft prompt. The main drawback of this approach was the lack of generalization that necessitated training for each lexically different instruction. To tackle the generalization problem, Mu et al. (2023) proposed to learn a Llama-7b to compress instruction to virtual tokens, but only compress instruction was not powerful enough since the demonstrations were much longer in practice. To compress the demonstrations, Chevalier et al. (2023) proposed AutoCompressor to recurrently generate compressed virtual tokens based on a fine-tuned Llama Zhang et al. (2022). However, AutoCompressor broke the independence of demonstrations, and the recurrent compression increased inference latency. Ge et al. (2023) proposed ICAE that employed a LoRA-adopted Llama-7b Touvron et al. (2023) to compress the processed demonstrations to compact virtual tokens, while ICAE still struggled to overcome quite long inputs.
@@ 2.2 Extractive Compression
Apart from employing soft prompts, researchers also endeavored to shorten prompts by extracting informative tokens from the original ones Li (2023); Jiang et al. (2023), namely token pruning Kim et al. (2022) or token merging Bolya et al. (2022). Recent works like LLMLingua Jiang et al. (2023) and Selective Context Li (2023) shared similarities but diverged on whether to eliminate tokens with high or low Perplexity (PPL). LLMLingua emphasized tokens with high PPL, attributing them as more influential, resulting in achieving outstanding performance. As mentioned in their paper, extractive compression methods encountered Out-of-Distribution (OoD) issues between the extractor and the target LLM. To reconcile this, they fine-tuned Alpaca-7b Taori et al. (2023) using the Alpaca dataset Taori et al. (2023) to perform the alignment. However, UniICL naturally bypassed the distribution-aligned module, since the compressor and the target LLM were the same.
Methods Additional Compressor Compression Tool # Trainable Parameters Train Size LLMLingua YES Pruning 7B 57k AutoCompressor NO Soft Prompt 7B UNKNOWN ICAE YES Soft Prompt 70M 240k UniICL NO Soft Prompt 17M 30k
Table 1: Comparsion among recent compression methods and UniICL. Compression Tool represents the involved compression technique of different methods.
@@ 3 Methodology
We propose UniICL, a parameter-efficient ICL framework that unifies demonstration compression, demonstration selection, and response generation via a single LLM. As for the selection of the underlying LLM, previous work has proved that the Decoder-only model performs better than the Encoder-Decoder model in prompt compression Mu et al. (2023). We follow this conclusion and employ popular Vicuna-7b Zheng et al. (2023) and BlueLM-7B [^2] We mainly discuss UniICL built on Vicuna-7B, the experiments of BlueLM will be exhibited in Appendiex B. Team (2023) as the underlying backbone in UniICL.
Figure 2: Differences of compression methods in formulating compressed virtual tokens in ICL.
We present a comparison of training costs between UniICL and other recent compression methods in Table 1. Additionally, we illustrate differences in formulating virtual tokens for compression methods based on the soft prompt in Figure 2.
To explain plainly, we ideally assume the compressor within three compression methods based on soft prompts has the window limitation of $L$, and has the same compression ratio, ignoring the length of soft prompts. In the 2-shot scenario, demonstrations $D_{1}$ and $D_{2}$ have a length of $L$. Considering the AutoCompressor, the concatenated demonstrations will be divided back into two segments, and the AutoCompressor compresses each segment step-by-step, bringing two times non-parallel compression. When it comes to ICAE, merely $D_{1}$ is accessible for the compressor and others will be read by no means. AutoCompressor shows advantages in the readable prompt length, but is short in efficiency, while ICAE has a constant compression complexity but struggles to approach relatively long inputs. Additionally, AutoCompressor makes the compression of later demonstrations be conditioned on the previous ones, which breaks demonstration independence in ICL.
Combining the advantages of AutoCompressor and ICAE, UniICL compresses demonstrations independently and introduces virtual tokens concatenation to overcome long prompt challenges. In the $N$-shot settings, the number of practical compression steps can be calculated as $\lceil\frac{N}{k}\rceil$, where $k$ indicates that a single GPU is capable of compressing $k$ demonstrations in a batch. When the GPU capacity is sufficient, k equals $N$, which is the scenario of ICAE that compresses all segments in a time but UniICL drops nothing, while it degenerates to the AutoCompressor scenario that compresses segments step-by-step, when the GPU capacity is only sufficient to set $k=1$.
Figure 3: Demonstration compression.
@@ 3.1 Demonstration Compression
UniICL introduces memory slots [M], a learnable embedding initialized from a rarely used embedding of the target LLM. UniICL activates the compression slots to absorb information from demonstrations in the forward propagation of frozen Vicuna, as illustrated in Figure 3. We first attach $k$ compression slots $=k\times\text{[M]}$M] to each demonstration $D_{i}$, formatting modified prompt fed to the Vicuna. Then, frozen Vicuna forwards the modified prompts and outputs the last hidden states $H_{i}=(h_{1},h_{2},...,h_{k})$ on top of the $k$ compression slots, dropping the others333$\oplus$ means token-level concatenation.:
_$\_,H^{i}=\mathrm{forward}(D_{i}\oplus M)$ (1)
Due to the attention mechanism, $H^{i}$ is compelled to attend to the preceding actual tokens. Then, UniICL inserts a projection layer to convert $H^{i}$ into LLM-acceptable compressed virtual tokens $C^{i}=(c^{i}_{1},c^{i}_{2},...,c^{i}_{k})$:
$c^{i}_{j}=W_{p}\cdot h^{i}_{j},$ (2)
@@ 3.2 Demonstration Selection
Except for substituting the origin demonstrations for generation, compressed virtual tokens $C^{i}$ are representative of demonstrations that can be again applied for demonstration selection, as illustrated in Figure 4. Specifically, given a query $Q$ and its candidate demonstrations $(D_{1},D_{2},...,D_{n})$, UniICL obtains their latent representation after average pooling:
$\bar{C^{i}}_{Q/D}=\frac{1}{k}\sum_{j=1}^{k}c_{j}.$ (3)
We define the $i$-th demonstration saliency score $S_{i}$ via the cosine similarity between Query and demonstrations:
$S_{i}=\mathrm{cos}(\bar{C_{Q}},\bar{C}_{D}^{i}).$ (4)
Demonstrations are then reranked according to their saliency scores.
3.3 In-context Generation
We employ the frozen Vicuna again to generate responses with the guiding of concatenated virtual demonstrations and queries, as illustrated in Figure 5. For $m$-shot in-context learning, we obtain $m$ virtual tokens after demonstration compression and selection, denoted as $C^{1}$ to $C^{m}$. Then, we horizontally concatenate them, keeping their relative position unmodified. Finally, the concatenated virtual tokens together with actual inference inputs are fed into Vicuna, performing auto-regressive generation as normal:
$y_{i}=\mathrm{generate}(C^{1},...,C^{m};Q;y_{<i})$ (5)
Except for the generative manner, virtual tokens are conveniently transferred to close-ended evaluation for understanding tasks through providing the candidate answers and measuring the label space PPL  [^4] https://huggingface.co/docs/transformers/perplexity, e.g. $(ppl^{+},ppl^{-}$) for sentiment classification:
$y=\mathrm{argmin}(ppl^{+},ppl^{-}),$ (6)
where answers with PPL closest to 1 are judged to be the current prediction.
Figure 5: In-context generation.
@@ 3.4 Training
The trainable parameters in UniICL are merely 17M originating from the projection layer $W_{p}$ and the introduced compression slot [M]. The projection layer is optimized with the language modeling objective of Vicuna to learn a base compression model. Then InfoNCE He et al. (2020) joint with language modeling objective are used to augment the demonstration selection ability of the base compression model:
@@
Specifically, we slice the source input of each training instance into two parts and randomly compress one, denoted the compression part as $x_{c}$ and the unmodified as $x_{u}$. Afterward, we attach $M$ to $x_{c}$ and get virtual tokens $C$ on top of the compression slots, as described in Equ. 1 and Equ. 2. Therefore, the language modeling loss $\mathcal{L}_{lm}$ is obtained as:
$\mathcal{L}_{lm}=-\frac{1}{|y|}\sum_{t=0}logP(y_{t}|Q;C;y_{<t}),$ (8)
where $y$ is the reference label of the current training instance. Additionally, to approach the large-shot settings without significant truncation, we introduce concatenation compression. When $x_{c}$ exceeds the window limitation for compression, UniICL further divides $x_{c}$ into acceptable ranges and compresses them independently to get local virtual tokens. Then, these virtual tokens of several segments will be concatenated to formulate global virtual tokens to replace $x_{c}$.
Consequently, we utilize contrastive learning for selection augmentation and mine positives and negatives as illustrated in Figure 6. Specifically, given each training instance $Q$ and $n$ candidate demonstrations $(D_{1},D_{2},...,D_{n})$ from two non-crossing training subsets, we employ Vicuna to calculate the PPL concerning the golden label of $Q$, denoted as $ppl^{Q}$. Then, we provide the $i$-th demonstration and calculate PPL concerning the golden label of $Q$, denoted as $(ppl^{D}_{i},i\in[1,n])$. We count $ppl^{Q}$ as the baseline and calculate candidate relative PPL gains:
@@
Hence, $C_{D}^{+}$ $(C_{D}^{-})$ is the representation of demonstrations $D^{+}$ ($D^{-}$) that furthest reduces (increases) $ppl^{Q}$ as processed in Equ. 3, and the contrastive loss $\mathcal{L}_{ctr}$ can be formulated as:
$\mathcal{L}_{ctr}=\frac{exp(cos(C_{Q},C_{D}^{+}))}{exp(cos(C_{Q},C_{D}^{+}))+exp(cos(C_{Q},C_{D}^{-}))}.$ (10)
@@
Figure 6: Negatives mining pipeline.

### v3

# UniICL: An Efficient Unified Framework Unifying Compression, Selection, and Generation

@@ Abstract
In-context learning (ICL) enhances the reasoning abilities of Large Language Models (LLMs) by prepending a few demonstrations. It motivates researchers to introduce more examples to provide additional contextual information for the generation. However, existing methods show a significant limitation due to the problem of excessive growth in context length, which causes a large hardware burden. In addition, shallow-relevant examples selected by off-the-shelf tools hinder LLMs from capturing useful contextual information for generation. In this paper, we propose UniICL, a novel Unified ICL framework that unifies demonstration compression, demonstration selection, and final response generation. Furthermore, to boost inference efficiency, we design a tailored compression strategy that allows UniICL to cache compression results into Demonstration Bank (DB), which avoids repeated compression of the same demonstration. Extensive out-of-domain evaluations prove the advantages of UniICL in both effectiveness and efficiency.
UniICL: An Efficient Unified Framework Unifying Compression, Selection, and Generation
[…]
Figure 1: (a) Prompt compression methods that indiscriminately compress both demonstrations and queries.(b) Retrieval-based demonstration selection methods select lexical demonstrations. (c) UniICL discriminately compresses demonstrations and performs selection upon the compression results.
@@ 1 Introduction
In-context learning (ICL) Brown et al. (2020); Xie et al. (2021); Wang et al. (2023b) to enhance the reasoning ability of Large Language Models (LLMs) with a few demonstrations prepended Wang et al. (2023d); Yang et al. (2023); Wei et al. (2023); Wang et al. (2023a); Min et al. (2022). Inspired by its outstanding performance, researchers explored applying ICL on many tasks such as text summarization Wang et al. (2023d); Yang et al. (2023); Gao et al. (2024a), sentiment classification, and linguistic acceptability Min et al. (2022); Wang et al. (2019). However, two challenges hinder the impact of ICL currently: (1) concatenated demonstrations directly surge the input length, causing a large hardware burden; (2) the prepended demonstrations are randomly sampled or selected via off-the-shelf tools which tend to provide shallow relevant demonstrations, hindering LLMs from capturing useful contextual information for generation. Existing work tackles the two challenges separately.
To alleviate input length surge, on the one hand, many efforts are made in modifying model architecture to accommodate longer contexts Zheng et al. (2022); Wu et al. (2022); Ding et al. (2023); Bulatov et al. (2023). These methods usually require training models from scratch, and models with a million context windows still struggle to overcome performance degradation Liu et al. (2024). On the other hand, recent studies attempt to shorten inputs through prompt compression Wingate et al. (2022); Mu et al. (2023); Jiang et al. (2023); Ge et al. (2023); Gao et al. (2024b). However, these compression methods are not applicable to ICL because they indiscriminately compress both demonstrations and queries into virtual tokens. For instance, as illustrated in Fig. 1(a), the task entails justifying whether the query is grammatically acceptable. The latter generator makes responses only according to virtual tokens generated by the compressor, resulting in a wrong answer [^1] I hope to would study in Facnce (France). More importantly, current compression methods are costly to train Wingate et al. (2022); Mu et al. (2023); Jiang et al. (2023), and compressors are either limited to compressing within the original model’s allowed input length Mu et al. (2023); Jiang et al. (2023); Ge et al. (2023) or bringing significant inference latency Wingate et al. (2022).
Retrieval-based In-context Example Selection (RICES) methods Alayrac et al. (2022) integrate an off-the-shelf pre-training model to select demonstrations similar to the queries at a shallow level. These demonstrations usually contain redundant information and bring minimal benefits for the final generation Liu et al. (2021); Ram et al. (2023); Wang et al. (2024). Existing work attempts to train the retrieval model and the generator in an end-to-end manner, which has shown better performance in in-domain datasets Wang et al. (2023c); Qiao et al. (2024). However, this approach still performs poorly in out-of-domain datasets. For instance, as shown in Fig. 1(b), the retriever selects an example lexically similar to queries but has contrasting labels. Then, the LLM is misled and responds with a wrong answer.
In light of challenges in ICL, we turn to leverage the inherent understanding ability of LLMs developed during pre-training. We accordingly propose a Unified ICL (UniICL) framework, which unifies demonstration compression, demonstration selection, and response generation. As shown in Fig. 1(c), for lightweight training, in UniICL, both the compressor and generator are initialized from the same LLM and kept frozen. An adapter is introduced to align the compressor with the generator, and [M] is a learnable embedding called Memory Slot which is attached behind demonstrations for compression. Therefore, UniICL only contains 17M trainable parameters. The LLM compressor first compresses each demonstration from the training set and queries into Memory Tokens independently on top of Memory Slots. Then, UniICL selects $n$ most relevant demonstrations based on the similarity of Memory Tokens between queries and demonstrations. Finally, Memory Tokens of selected demonstrations are concatenated to formulate a global in-context sequence, together with queries fed into the generator for response generation. Due to independent compression, the compressor gets rid of the input window limitation of original LLMs as the number of demonstrations increases. In addition to improvements in window limitation, the tailored compression strategy further makes improvements to ICL efficiency. Specifically, UniICL caches Memory Tokens of different demonstrations to configure the Demonstration Bank (DB) for future reusing as shown in Fig. 2. Therefore, repeated compression of the same demonstration is not necessary, which significantly boosts model efficiency in Fig. 8. Extensive out-of-domain evaluation indicates UniICL achieves substantial improvements compared with other baselines. Our main contributions are as follows:
Figure 2: The workflow of Demonstration Bank.
  • To our knowledge, we are the first to propose a unified ICL framework with 17M trainable parameters.
@@
  • UniICL proposes configuring the Demonstration Bank to avoid repeated compression for the same demonstration, which significantly boosts ICL efficiency.
  • Different from the indiscriminate compression of previous studies, UniICL proposes a tailored compression strategy for ICL, achieving substantial improvements compared with other baselines.
@@ 2.1 Soft Prompt Compression
Recently, researchers attempted to utilize soft prompts to convert actual tokens to dense-information virtual tokens. Mostly from a distillation perspective, Wingate et al. (2022) aligned the teacher model and the student model, where the teacher model accepted the actual task instruction while the student model fed the soft prompt. The main drawback of this approach was the lack of generalization that necessitated training for each lexically different instruction. To tackle the generalization problem, Mu et al. (2023) proposed to learn a Llama-7b to compress instructions to virtual tokens, but only compressing instructions was not powerful enough since the demonstrations were much longer in practice. To compress longer prompts, Chevalier et al. (2023) proposed AutoCompressor to recurrently generate compressed virtual tokens based on a fine-tuned Llama Zhang et al. (2022). However, AutoCompressor broke the independence of demonstrations, and the recurrent compression increased inference latency. Ge et al. (2023) proposed ICAE that employed a LoRA-adopted Llama-7b Touvron et al. (2023) to compress the processed demonstrations to compact virtual tokens, while ICAE still struggled to overcome quite long inputs.
@@ 2.2 Extractive Compression
Apart from employing soft prompts, researchers also endeavored to shorten prompts by extracting informative tokens from the original ones Li (2023); Jiang et al. (2023), namely, token pruning Kim et al. (2022) or token merging Bolya et al. (2022). Recent works like LLMLingua Jiang et al. (2023) and Selective Context Li (2023) shared similarities but diverged on whether to eliminate tokens with high or low Perplexity (PPL). LLMLingua emphasized tokens with high PPL, attributing them as more influential, resulting in achieving outstanding performance. As mentioned in their paper, extractive compression methods encountered Out-of-Distribution (OOD) issues between the extractor and the target LLM. To reconcile this, they fine-tuned Alpaca-7b Taori et al. (2023) using the Alpaca dataset Taori et al. (2023) to perform the alignment.
@@ 3 Methodology
Previous compression methods are not tailored for ICL, and they are either bound by serious inference latency or poor performance, as demonstrated in Appendix A. We propose UniICL, a unified ICL framework that unifies demonstration compression, demonstration selection, and response generation. As for the selection of the underlying LLM, previous work has proved that the Decoder-only model performs better than the Encoder-Decoder model in prompt compression Mu et al. (2023). We follow this conclusion and adopt Vicuna-7B Zheng et al. (2023) as the underlying backbone in UniICL.
Figure 3: Demonstration compression. $k$ Memory Slots are attached behind each demonstration.
@@ 3.1 Demonstration Compression
[…]
$H^{i}=f_{\theta}(D_{i}^{L_{i}\times d}\oplus M^{k\times d}),$ (1)
where $L_{i}$ is the $i-$th demonstration length, $d$ is the embedding dimension and $\oplus$ means token-level concatenation. Due to the attention mechanism, $H^{i}$ is compelled to attend to the preceding actual tokens. Then, UniICL applies a linear layer as the adapter for efficiency to convert $H^{i}$ to Memory Tokens $C^{i}=(c^{i}_{1},c^{i}_{2},...,c^{i}_{k})$, performing alignment between the compressor and the generator [^2] Linear layer is enough for UniICL as features have interacted with each other during compression.:
$c^{i}_{j}=W_{p}^{d\times d}\cdot h^{i}_{j},$ (2)
@@ 3.2 Demonstration Selection
Memory Tokens $C^{i}$ naturally summarize the demonstrations in latent space, and UniICL performs demonstration selection based on the similarity between queries and demonstrations as shown in Fig. 4. Specifically, given a query $Q$ and its candidate demonstrations $(D_{1},D_{2},...,D_{n})$, UniICL obtains their representations used for selection by average pooling $C_{\{Q,D\}}$:
$\bar{C^{i}}_{\{Q,D\}}=\frac{1}{k}\sum_{j=1}^{k}c_{j}.$ (3)
We define the $i$-th demonstration saliency score $S_{i}$ as the cosine similarity between $\bar{C_{Q}}$ and $\bar{D_{i}}$:
$S_{i}=\mathrm{cosine\_similarity}(\bar{C_{Q}},\bar{C}_{D}^{i}).$ (4)
3.3 Generation
We employ the frozen Vicuna again to generate responses with the guidance of concatenated Memory Tokens and queries, as illustrated in Fig. 5. For $m$-shot in-context learning, we obtain $m$ spans of Memory Tokens after demonstration compression and selection, denoted as $C^{1}$ to $C^{m}$. Then, we horizontally concatenate them, keeping their relative position unmodified. Finally, the concatenated Memory Tokens together with actual queries are fed into Vicuna, performing auto-regressive generation $g_{\theta}$ as normal:
$y_{i}=g_{\theta}(C^{1},...,C^{m};Q;y_{<i}).$ (5)
Except for the generative manner, Memory Tokens apply close-ended evaluation for understanding tasks as normal through measuring the perplexity of candidate choices  [^3] https://huggingface.co/docs/transformers/perplexity.
Figure 5: In-context generation. The Memory Tokens from different demonstrations are concatenated horizontally at the input end of Vicuna.
@@ 3.4 Training
The trainable parameters in UniICL are merely 17M originating from the projection layer $W_{p}$ and the introduced Memory Slot [M]. The linear layer is optimized with the language modeling objective $\mathcal{L}_{lm}$ of Vicuna to learn a base compression model. Then InfoNCE He et al. (2020) joint with language modeling objective are used to augment the demonstration selection ability of the base compression model:
@@
Specifically, we slice the source input of each training instance into two parts and randomly compress one. The compressed part is denoted as $x_{c}$ and the uncompressed part is denoted as $x_{u}$. Afterward, we attach the Memory Slot sequence $M$ behind $x_{c}$ and get Memory Tokens $C$ on top of the Memory Slots, as described in Eq. 1 and Eq. 2. Therefore, the language modeling loss $\mathcal{L}_{lm}$ is obtained as:
$\mathcal{L}_{lm}=-\frac{1}{|y|}\sum_{t=0}logP(y_{t}|x_{u};C;y_{<t}),$ (7)
where $y$ is the reference label of the current training instance. Additionally, to approach the large-shot settings without significant truncation, we introduce concatenation compression. When $x_{c}$ exceeds the window limitation for compression, UniICL further divides $x_{c}$ into acceptable ranges and compresses them independently to get local Memory Tokens. Then, these Memory Tokens from different segments will be concatenated to formulate global virtual tokens to replace $x_{c}$, applying Eq. 7 to optimize models as well.
We obtained a base compression model that has learned to compress and understand concatenated Memory Tokens after the first-phase training mentioned. Subsequently, we utilize contrastive learning for selection augmentation and mine positives and negatives as illustrated in Fig. 6. Specifically, given each training instance $Q$ and $n$ candidate demonstrations $(D_{1},D_{2},...,D_{n})$ from two non-crossing training subsets, we employ Vicuna to calculate the PPL concerning the golden label of $Q$, denoted as $ppl^{Q}$ to find useful demonstrations for generation. Then, we provide the $i$-th demonstration and calculate PPL concerning the golden label of $Q$, denoted as $(ppl^{D}_{i},i\in[1,n])$. We count $ppl^{Q}$ as the baseline and calculate candidate relative PPL gains:
@@
After finding demonstrations $D^{+}$ ($D^{-}$) that furthest reduces (increases) $ppl^{Q}$, we obtain their representation $C_{D}^{+}$ $(C_{D}^{-})$ as processed in Eq. 3. The contrastive loss $\mathcal{L}_{ctr}$ can be formulated as:
$\mathcal{L}_{ctr}=\frac{\exp(cos(C_{Q},C_{D}^{+}))}{\exp(cos(C_{Q},C_{D}^{+}))+\exp(cos(C_{Q},C_{D}^{-}))}.$ (9)
@@
Figure 6: Contrastive examples mining pipeline. Finds demonstrations benefit/hinder the final generation according to the PPL.